determine the odds for and against each event for the given probabliity:
P(H)= 1/4
P(G)= .1

Answer by sudhanshu_kmr(1152)


P(H) = 1/41 , P'(H) = 40/41
odds in favor = 1/40 , odds in against = 40/1


P(G) = .1 = 1/10 P'(G)= 9/10
odds in favor = 1/9 , odds in against = 9/1 = 9